OpenStudy (anonymous):

Find two numbers to make the equation true. 3m2 = 147 @Directrix

Directrix (directrix):

3m2 = 147 Divide both sides by 3

Directrix (directrix):

m² = 147/3

Directrix (directrix):

What is 147 divided by 3

Directrix (directrix):

@desfields901

OpenStudy (anonymous):

58

Directrix (directrix):

No, try again.

OpenStudy (anonymous):

my computer is a bit older so its losing signal and crashing so i have to keep refreshing an it takes a while sometimes

OpenStudy (anonymous):

49 sorry i did 173/3

Directrix (directrix):

m^2 = 49 That means what number times itself = 49. What is one number that behaves that way? You may have a square root key on your calculator. I don't know but this square root is pretty easy to get out of your head.

Directrix (directrix):

I am not typing. This crazy message says that but I am not.

Directrix (directrix):

Go to Google and in the search box, enter this: "square root of 49" and tell me what it says.

Directrix (directrix):

@desfields901

OpenStudy (anonymous):

it works

OpenStudy (anonymous):

7

Directrix (directrix):

Listen up. Every number has two square roots. 7 is one of the square roots of m. -7 is the other. So, remember the calculator will give you the positive root but you have to get the negative root out of your head.

Directrix (directrix):

I mean to say square roots of 49, not m.

OpenStudy (anonymous):

okay

Directrix (directrix):

3 times (7)^2 = 147 AND 3 times (-7)^2 = 147

Directrix (directrix):

Next Practice Question

OpenStudy (anonymous):

okay so how did we get 7 again

Directrix (directrix):

Divide 147 by 3 to get x^2 = 49. Then take square roots to get 7 and -7. We divided by 3 to begin with because 3x^2 = 147 and you have to divide by 3 to isolate the x^2 term and then solve for x.

Directrix (directrix):

you got the 7 by using Google, entering square root of 49

OpenStudy (anonymous):

7,-7

Directrix (directrix):

Yes. Look. Click on this link. http://openstudy.com/users/desfields901# When you get there, click on "quetions asked" Find this question back there and come in the back door. It is quicker that way when 400+ people are on the site. We'll just stay in one thread once we get there and you can pay me later. We need to pick up some speed.

Directrix (directrix):

You can't post questions from back there so we'll stay in the thread and you can post questions in that thread and not get hung up at the front door of this site.

OpenStudy (anonymous):

what does that mean

Directrix (directrix):

If you would click on this. link, I could show you. http://openstudy.com/users/desfields901# But if you don't want to, don't/